OpenStudy (anonymous):

Set up and solve a proportion A ski course falls 50 feet in every 500 feet of horizontal run. If the total horizontal run is 1/2 mile, find the height of the hill. (Hint: 5,280 feet = 1 mile.) _________ft.

OpenStudy (lilg132):

5280 / 2 = 2640 2640 / 500 = 5.28 5.28 * 50 = 264 feet has fell 2640 - 264 = 2376 feet is the size of the hill

OpenStudy (lilg132):

*height of the hill i meant

OpenStudy (lilg132):

dont forget to click good answer :)